# Efficient Cloud-Edge Collaborative Inference for Object Re-Identification

### v1

@@ Abstract
Current object re-identification (ReID) system follows the centralized processing paradigm, i.e., all computations are conducted in the cloud server and edge devices are only used to capture and send images. As the number of videos experiences a rapid escalation, this paradigm has become impractical due to the finite computational resources. In such a scenario, the ReID system should be converted to fit in the cloud-edge collaborative processing paradigm, which is crucial to boost the scalability and practicality of ReID systems. However, current relevant work lacks research on this issue, making it challenging for ReID methods to be adapted effectively. Therefore, we pioneer a cloud-edge collaborative inference framework for ReID systems and particularly propose a distribution-aware correlation modeling network (DaCM) to make the desired image return to the cloud server as soon as possible via learning to model the spatial-temporal correlations among instances. DaCM embeds the spatial-temporal correlations implicitly included in the timestamps into a graph structure, and it can be applied in the cloud to regulate the size of the upload window and on the edge device to adjust the sequence of images, respectively. Traditional ReID methods can be combined with DaCM seamlessly, enabling their application within our proposed edge-cloud collaborative framework. Extensive experiments demonstrate that our method obviously reduces transmission overhead and significantly improves performance. We will release our code and model.
@@ 1 Introduction
Figure 1: The overview of our approach. The DaCM Network is performed in both cloud server and edge devices for adjusting uploading batch size $b^{i}$ and image order in the uploading sequence. The red solid denotes the data flow provided by our method, and the gray dashed line denotes the previous data flow without our method.
@@
The object re-identification (ReID) system facilitates users in searching objects accurately across diverse scenes and views, significantly alleviating manual overhead. Serving as a fundamental task in visual surveillance, an increasing number of innovative technologies are being introduced, including elaborate feature extractors [25, 14, 11, 19], data transmission schemes [16], and inference strategies [41], which greatly promote the performance. Consequently, the ReID system has been widely applied to real-world scenarios, playing an important role in traffic monitoring, safety management, etc.
Due to its inherent cross-scene characteristics, the current ReID system mainly comprises one cloud server, multiple edge devices, and a connected network for communication, adhering to centralized processing patterns. The edge devices only capture images and upload them to the cloud server through the connected network, while the cloud server receives them and employs a ReID method for feature extraction and similarity calculation. Ultimately, images are returned to the user in accordance with the resulting similarity. However, with the rapid growth in the number of cameras, such processing patterns put undue pressure on the network bandwidth and computing/storage capacity of the cloud server, which results in considerable service delays and a diminished user experience. Therefore, in line with the current development trend [2], the ReID system should integrate into a cloud-based collaborative framework.
@@
Previous cloud-edge collaborative visual methods primarily concentrate on the training phase. For instance, FedReID [42] and FedUReID [43] embed the federated learning into ReID system, delving into strategies to exploit distributed data to continuously optimize the feature extractor, thereby enhancing search accuracy. Besides, some works [17] proposes to deploy a segment of the deep neural network to the edge devices, mitigating the computing cost of the cloud server, although it still necessitates a substantial amount of data transmission. We argue that the inference phase holds greater significance for a practical ReID system, and a meticulous scheme should be developed to fully leverage the advantages of both cloud server and edge devices. Therefore, in this paper, in contrast to previous methods, we introduce a pioneering cloud-edge collaborative ReID system, which places a heightened emphasis on optimizing the efficiency of the inference process, a domain that has been underexplored in existing research.
@@
As shown in Fig. 1, in the cloud-edge collaborative framework, one query image is dispatched to each edge device, where it extracts its feature using a local visual backbone and compares this feature with all local gallery images. The resulting similarity sorts the local gallery images to create the uploading sequence. Due to transmission limitations, There is an upper limit to the amount of data a cloud server can accept at one time, so images are uploaded in batches. To achieve an efficient inference, the user’s desired image should be returned to the cloud server promptly. Therefore, two key points in our framework are: (i) the edge device with the desired image should have a higher chance of transmitting the image to the cloud server, and (ii) the desired image should be positioned at the beginning of the upload sequence.
In addition to images, we find that the timestamp is another free launch from the ReID system, which can be effortlessly obtained and utilized to improve the inference. Therefore, to achieve the above goals, we specifically introduce a distribution-aware correlation modeling network (DaCM), transforming learning spatial-temporal correlation into a classification problem. Initially, it embeds spatial-temporal information into a graph structure, followed by several graph convolutional layers to propagate information from to all nodes representing cameras. Finally, we predict the probability that the target will appear on each camera according to the produced node features. Having learned the camera network topology, DaCM can be deployed on the cloud server to allocate transmit chances for each device, and also be deployed on the edge device to re-rank the image sequence. Promisingly, DaCM also enables the system to perform time-constrained ReID (tcReID), i.e. given a target time, the system can return the temporally closest image of the same identity, which is a more challenging task but not effectively addressed by previous methods.
Furthermore, since we propose a new ReID paradigm (cloud-edge collaborative inference) and task (tcReID), traditional evaluation protocols do not fully showcase its capabilities. Thus, we propose several new evaluation protocols including mean Transmission Number (mTN), Precise Rank@K (pR-K), and mean Precise Rank (mpR). Extensive experimental results demonstrate that our method not only significantly improves traditional evaluation performance but also enhances bandwidth utilization, and reduces searching time, showcasing its superiority. The contributions of our work can be summarised as:
  • We propose a cloud-edge collaborative inference framework for ReID system, which can fully use both the advantages of cloud and edges for efficient inference.
@@
  • We design a distribution-aware correlation modeling network to model the spatial-temporal correlations, which boosts the efficiency of the system via enlarging the return probability of the desired image.
@@
  • We introduce several evaluation protocols and conduct extensive experiments on various benchmark datasets to demonstrate the superiority of our method.
@@ 2.1 Object Re-identification
Object re-identification (ReID) is one of the most important tasks in visual surveillance. It aims at associating person/vehicle images with the same identity captured by different cameras in diverse scenarios. Earlier methods are type-specific, relying on specific attributes of the object, and are applicable only to a particular type of objects, such person ReID [38, 1, 4, 39] and vehicle ReID [23, 24, 22]. As methods continue to advance, there is a growing trend towards developing generic ReID methods that are agnostic to the type of object being applied. The focus is on creating general deep learning technologies, including feature extractors [25, 14, 11, 19], metric strategies [36], and loss functions [4, 31]. FastReID [10] integrates the majority of the above techniques, and a simple ResNet50 [9] backbone can deliver outstanding performance on both person ReID and vehicle ReID datasets. He et al. [11] propose a model named TransReID, which adopts the Vision Transformer [7] as the backbone and achieves superior performance for both person and vehicle ReID tasks. All the above methods can be employed in our cloud-edge collaborative framework, partnering with DaCM for efficient inference.
Since spatial-temporal information can be painlessly obtained for ReID data, some ReID methods [13, 5, 33] incorporate it to filter out unreasonable samples. Compared with them, our approach has several obvious differences : (i) Previous methods generate the spatial-temporal distribution through frequency statistics, whereas our approach employs a deep neural network to learn such correlations; (ii) Previous methods still adhere to centralized patterns, whereas our approach is developed within a cloud-edge collaborative framework; (iii) Previous methods only use such information to enhance performance, whereas our approach improves performance while achieving efficient reasoning.
@@ 2.2 Cloud-Edge Collaborative Methods
Emerging cloud-edge collaboration approaches showcase their superiority in various systems and communication technologies. Noteworthy instances of these advanced methodologies are evident in seminal works, such as the collaborative occluded face recognition architecture [37], the open-source framework SmartEye for real-time video analytic [34], and the video service enhancement within an edge-cloud collaboration framework [35]. The adaptation of cloud-device collaboration sensitive to changing environments [8], the real-time surveillance video analysis in Cloud-Edge architecture [12], and the Classification Driven Compression framework for reducing deep learning bandwidth consumption [6] further underscore the versatility and impact of these collaborative approaches. Different from these methods, we focus on the ReID task and aim to achieve efficient inference but not model optimization. In particular, we design a cloud-edge collaborative ReID system, in which a DaCM is performed to push the desired image return to the cloud server promptly via modeling the spatial-temporal correlations.
@@ 3 Problem Definition
We initially present the problem definition of efficient cloud-edge collaborative inference by outlining the framework’s pipeline and introducing the tasks of ReID.
Given a query image $\mathcal{I}^{q}$ and its spatial-temporal information $\{l^{q},t^{q},c^{q}\}$ ($l^{q}$, $t^{q}$, and $c^{q}$ denote the identity ID, taken timestamp, and camera ID, respectively), the ReID system sends these information to every edge device. Then, for the $i$-th edge device, it extracts the deep feature $\mathbf{f}^{q}$ from $\mathcal{I}^{q}$ via the local network and compute the similarity $\mathbf{S}\in\mathbb{R}^{N_{i}}$ between $\mathbf{f}^{q}$ and the features $\mathbf{G}^{i}\in\mathbb{R}^{N_{i}\times E}$ of all $N_{i}$ gallery images $\mathcal{G}^{i}$, where $E$ is the dimension number of deep features. In conventional methods, the uploading sequence is generated depending solely on $\mathbf{S}$, but in our approach, we assign the index of each image via considering jointly the visual similarity and spatial-temporal correlations, which can for formulated as:
@@
where $d$ denotes the distance function for visual features, $e$ denotes the function of computing spatial-temporal correlations, and $\theta$ is used to combine them together.
Then, according to $\mathbf{s}^{i}\in\mathbb{R}^{N_{i}}$, the uploading image sequence is produced as $\mathcal{U}^{i}=\{\ldots,\mathcal{G}^{i}_{p},\ldots,\mathcal{G}^{i}_{q}\ldots\}$, which subjects to $0\leq p,q\leq N_{i}$, and $s^{i}_{p}<=s^{i}_{q}$. As described in the introduction, these images will be uploaded to the cloud server in batches. If the $i$-th edge device can upload $b^{i}$ data each time, the transmission number can be formulated as:
[…]
$\epsilon$ is an index function that returns the index of the desired image in $\mathcal{U}^{i}$, whose implementation is contingent upon user requirements. For example, in the normal ReID task, it is acceptable for the system to just return one image with ID $l^{q}$, so $\epsilon$ represents the index of the first image in $\mathcal{U}^{i}$ with ID $l^{q}$. Note that the procedure is utilized to generate $TN$, showcasing the efficiency of inference. In a real-world scenario, the gallery image lacks a label, and it needs the user to verify the returned images. Finally, for Eq. (2), we select the minimum transmission number of all edge devices as the final $\mathrm{TN}$. From Eq.(1) and Eq.(2), it is evident that to achieve efficient inference (i.e., minimizing the value of $\mathrm{TN}$), $s^{i}_{j}$ of the desired image should be small, and $b^{i}$ should be large for the edge device containing the desired image. Therefore, the problem can be formulated as reducing $\mathrm{TN}$ by learning appropriate $s^{i}_{j}$ and $b^{i}$ from the spatial-temporal information.
In a ReID system, users usually not only seek returned images with the same ID as the query image but also aim to acquire images captured around a specific time, a scenario referred to as time-constrained ReID. So, if the images of one certain target are taken at different times, the sample space of the correct image is narrowed down to a single element. Thus, tcReID is a more challenging variant of ReID, and upon removing the time constraint, tcReID reverts to a normal ReID task.
@@ 4 Proposed Approach
In this section, we first describe how the proposed cloud-edge collaborative ReID system works, and how the DaCM network performs within the system to boost its efficiency.
@@
4.1 System Overview
Fig. 1 provides an overview of our proposed cloud-edge collaborative ReID system. The user inputs the request to the cloud server, and the query data is sent to all edge devices along with the bandwidth information $b^{i}$ produced by the DaCM network in the cloud server. For each device, it employs a receiver to accept the query information and then utilizes a visual backbone to extract deep features of the image. Additionally, the DaCM network takes in the spatial-temporal information of the query image ($t^{q},c^{q}$) and gallery data ($\mathbf{t}^{i}$), and outputs the spatial-temporal correlation, which is fused with the visual similarity to form the joint similarity. The gallery images are sorted based on the joint similarity to construct the uploading sequence. These images are sent to the cloud server in batches, the size of which is determined by $b^{i}$.
@@ 4.2 DaCM Architecture
The DaCM network performs an important role in the cloud-edge collaborative ReID system, and in this part we describe its architecture. As shown in Fig. 2, DaCM mainly consists of three components, a spatial-temporal embedding module, multiple Correlation Modeling (CoMo) blocks, and a final classifier.
Spatial-temporal embedding. Inspired by the positional encoding manner used in [32], we adopt the sinusoidal embedding to encode the temporal information. Denoted the query and target times as $t^{q}$ and $t^{d}$, we encode their difference in the formulation of:
$\mathbf{e}_{2i}$ $=\sin\left((t^{d}-t^{q})/\lambda^{2i/D}\right),$ (3)
$\mathbf{e}_{2i+1}$ $=\cos\left((t^{d}-t^{q})/\lambda^{2i/D}\right),$
where $i$ is the dimension index, $\mathbf{e}\in\mathbb{R}^{D}$ is the results embedding, and $D$ is the dimension of the embedding. $\lambda$ denotes the max period of the sinusoidal function. The wavelengths form a geometric progression from $2\pi$ to $10000\cdot 2\pi$.
For the spatial information, considering the fixed localizations of different cameras, we suggest employing learnable parameters $\mathbf{W}\in\mathbb{R}^{C\times D\times C\times D}$ to represent the spatial embedding. $C$ is the number of cameras and the nodes in the graph. When $c^{q}$ is provided as input, we choose $\mathbf{W}_{c^{q}}$ as the resulting spatial embedding. This is then combined with $\mathbf{e}$ to generate a graph structure:
$\mathbf{A}=\frac{1}{\sum_{j}^{D}e_{j}}\sum_{j}^{D}\mathbf{W}_{c^{q},j}*e_{j}+\mathbf{b}\in\mathbb{R}^{C\times D}.$ (4)
Correlation modeling block. As for CoMo blocks, they are used to propagate the information among nodes, and as shown in Fig. 2, one CoMo block is comprised of a Layer Normalization, a GCN layer, and a GeLU activation, which can be formulated as:
@@
where $\mathbf{E}^{l}\in\mathbb{R}^{C\times C}$ is the adjacency matrix, which can be learned during training, and $\mathbf{W}^{l}\in\mathbb{R}^{C\times C}$ denotes the transfer weight to update the information for each node.
@@
The MLP consists of the sequence of BatchNorm [15], ReLU, and fully connected layers.

### v2

# Towards Efficient Object Re-Identification with A Novel Cloud-Edge Collaborative Framework

@@ Abstract
Object re-identification (ReID) is committed to searching for objects of the same identity across cameras, and its real-world deployment is gradually increasing. Current ReID methods assume that the deployed system follows the centralized processing paradigm, i.e., all computations are conducted in the cloud server and edge devices are only used to capture images. As the number of videos experiences a rapid escalation, this paradigm has become impractical due to the finite computational resources in the cloud server. Therefore, the ReID system should be converted to fit in the cloud-edge collaborative processing paradigm, which is crucial to boost its scalability and practicality. However, current works lack relevant research on this important specific issue, making it difficult to adapt them into a cloud-edge framework effectively. In this paper, we propose a cloud-edge collaborative inference framework for ReID systems, aiming to expedite the return of the desired image captured by the camera to the cloud server by learning the spatial-temporal correlations among objects. In the system, a Distribution-aware Correlation Modeling network (DaCM) is particularly proposed to embed the spatial-temporal correlations of the camera network implicitly into a graph structure, and it can be applied 1) in the cloud to regulate the size of the upload window and 2) on the edge device to adjust the sequence of images, respectively. Notably, the proposed DaCM can be seamlessly combined with traditional ReID methods, enabling their application within our proposed edge-cloud collaborative framework. Extensive experiments demonstrate that our method obviously reduces transmission overhead and significantly improves performance.
Codehttps://github.com/bupt-wcm/AAAI-DaCM.git
@@ 1 Introduction
Object re-identification (ReID) (He et al. 2021, 2023; Li, Sun, and Li 2023; Luo et al. 2019; Huynh 2021; Ge et al. 2024; Chen et al. 2021; Fan et al. 2018; Fu et al. 2024; Qi et al. 2021; Qi, Wang, and Li 2017; Liu et al. 2018) aims to retrieve specific objects captured by non-overlapping cameras, which usually serves as a fundamental task in the field of multimedia processing. It can facilitate users in searching objects accurately across diverse scenes and views, significantly alleviating manual overhead in visual surveillance. With the increasing demands, ReID systems have been widely deployed in various real-world scenarios for vehicle or person searching, playing an important role in traffic monitoring, safety management, etc. Therefore, an expanding number of innovative technologies have been introduced to promote the accuracy of ReID system continuously, including establishing elaborate feature extractors (Luo et al. 2019; Huynh 2021; He et al. 2021; Li, Sun, and Li 2023), developing data transmission schemes (Jain et al. 2020), and designing inference strategies (Zhong et al. 2017).
Figure 1: Illustration of the difference between centralized and cloud-edge collaborative patterns for ReID systems.
Due to its intrinsic cross-scene nature, a ReID system typically consists of a central cloud server, multiple edge devices (such as cameras), and a communication network for transmitting images and associated data. Previous ReID methods typically follow a centralized processing pattern, where, as illustrated on the left of Fig. 1, edge devices merely capture images and upload all images to the cloud server via the connected network. The cloud server then utilizes a deep neural network to extract features and compute similarities between the query and returned images. However, with the rapid proliferation of cameras, this processing pattern imposes excessive strain on the communication network’s bandwidth and the cloud server’s computing and storage capacities, leading to significant service delays and a compromised user experience. Consequently, in line with current technological trends and driven by the advancement of device computing power (Angel et al. 2022), as depicted on the right of Fig. 1, the ReID system should seamlessly integrate into a cloud-based collaborative framework. Feature extraction should occur locally at the edge device, with partial images being uploaded to the cloud server based on the ReID model’s outputs, thereby alleviating the burden on network communication and cloud computing.
@@
Some previous methods (Zhuang et al. 2020; Zhuang, Wen, and Zhang 2021; Jiang et al. 2023) want to establish new-style cloud-edge collaborative frameworks, but they primarily concentrate on the training phase. For instance, FedReID (Zhuang et al. 2020) and FedUReID (Zhuang, Wen, and Zhang 2021) embed the federated learning into ReID system, delving into strategies to exploit distributed data to continuously optimize the feature extractor, thereby enhancing search accuracy. Besides, some works (Jiang et al. 2023) propose to deploy a segment of the deep neural network to the edge devices, mitigating the computing cost of the cloud server, although it still necessitates a substantial amount of data transmission. We argue that the inference phase also holds greater significance for a practical ReID system, and a meticulous scheme should be developed to fully leverage the advantages of both the cloud server and edge devices. Therefore, in contrast to previous methods, we introduce a pioneering cloud-edge collaborative ReID system that places a heightened emphasis on optimizing the efficiency and effectiveness of the inference process, a domain that has been under-explored in existing research.
@@
For a basic cloud-edge collaborative inference pipeline of the ReID system, when the user requests to search one certain object, one query image and its auxiliary information (denoted as query in Fig. 1) are initially dispatched to each edge device from the cloud server by a Transmitter. Then, the edge device extracts its feature via a local visual backbone and compares this feature with all local gallery images, and the resulting similarity is used to create the uploading sequence (denoted as sequence in Fig. 1). Due to transmission limitations, there is an upper bound to the number of data the cloud server can accept at one time, so the uploading sequence of images is uploaded in batches. The user checks the sequence and terminates this process when its desired image is returned. Therefore, to achieve an efficient and effective inference, the user’s desired image should be returned to the cloud server swiftly. Therefore, two key points in our framework are: (i) the edge device with the desired image should have a higher chance of uploading the image to the cloud server, and (ii) the desired image should be positioned at the beginning of the upload sequence.
To handle the points above, we specifically introduce a distribution-aware correlation modeling network (DaCM), which is deployed in both the cloud server to adjust the bandwidths of edge devices and each edge device to re-rank the image indexes in the uploading sequence. The input of DaCM is spatial-temporal data, i.e. the timestamps and camera ID of images, which can be effortlessly obtained from the ReID system. Initially, it embeds spatial-temporal correlations into a graph structure by learning such a problem: what is the likelihood that an object will appear again in camera $j$ after a time delay $t$, from where it was previously observed in camera $i$. After training, the topology of the camera network and the movement rules of the object in the current scene will be embedded into DaCM implicitly, so as to support the adjustment of the bandwidth allocated to the edge devices and the index of the image.
Furthermore, since we focus on a new ReID inference pattern, traditional evaluation protocols do not fully showcase the capabilities of proposed method. Thus, we propose several new evaluation protocols and their details will be described later. Finally, extensive experimental results demonstrate that our method improve the performance with a significant enhancement in accuracy and efficiency.
The contributions of our work can be summarised as:
  • To handle the rapidly growing number of videos, we propose an inference framework for ReID systems, which can evolve current methods into a cloud-edge collaborative pattern, enhancing both efficiency and effectiveness.
@@
  • To boost the system’s performance by increasing the return probability of the desired image, we design a Distribution-aware Correlation Modeling network that captures the spatial-temporal correlations of the scene.
@@
  • To demonstrate the superiority of our method, we introduce several evaluation protocols and conduct extensive experiments, with the results showcasing the significant enhancement achieved by our proposed framework.
@@ 2.1 Object Re-identification
Earlier ReID methods are type-specific, relying on specific attributes of the object, and are applicable only to a particular type of objects, such as person ReID (Zheng et al. 2015; Ahmed, Jones, and Marks 2015; Cheng et al. 2016; Zheng, Zheng, and Yang 2017) and vehicle ReID (Liu et al. 2016a, b, 2017). As methods continue to advance, there is a growing trend towards developing generic ReID methods (Luo et al. 2019; Huynh 2021; He et al. 2021; Li, Sun, and Li 2023; Ye et al. 2022; Cheng et al. 2016; Sun et al. 2020; He et al. 2021) that are agnostic to the type of object being applied. All the above methods can be employed in our cloud-edge collaborative framework, partnering with DaCM for efficient and effective inference.
Since spatial-temporal information can be effortlessly obtained in a ReID system, some ReID methods (Huang et al. 2016; Cho et al. 2019; Wang et al. 2019) incorporate it to filter out unreasonable samples. Compared with them, our approach has several obvious differences : (i) Previous methods generate the spatial-temporal distribution through frequency statistics, whereas our approach employs a deep neural network to learn such correlations; (ii) Previous methods still adhere to centralized patterns, whereas our approach is developed within a cloud-edge collaborative framework; (iii) Previous methods only use such information to enhance performance, whereas our approach improves performance while achieving efficient inference. As a similar work, Jain et al. also interpolate such information in object searching, but the proposed Spatula directly filters out many candidate images that leads that (i) the desired image may not be found even with the replay strategy, and (ii) it is hard to combined with neural networks for promising performance.
@@ 2.2 Cloud-Edge Collaborative Methods
Emerging cloud-edge collaboration approaches showcase their superiority in various systems and communication technologies. Noteworthy instances of these advanced methodologies are evident in seminal works, such as the collaborative occluded face recognition architecture (Zhang et al. 2023), the open-source framework SmartEye for real-time video analytic (Wang and Gao 2021), and the video service enhancement within an edge-cloud collaboration framework (Wu et al. 2021). The adaptation of cloud-device collaboration sensitive to changing environments (Gan et al. 2023), the real-time surveillance video analysis in Cloud-Edge architecture (Hou and Zhang 2021), and the Classification Driven Compression framework for reducing deep learning bandwidth consumption (Dong et al. 2020) further underscore the versatility and impact of these collaborative approaches. Unlike these methods, we focus on the ReID task and aim to achieve efficient and effective inference instead of model optimization.
@@ 3 Problem Definition
As shown in Fig. 2, given a query image $\mathcal{I}^{q}$ and auxiliary information $\{t^{q},c^{q},t^{d}\}$ ($t^{q}$ and $c^{q}$ denote the timestamp and camera ID of $\mathcal{I}^{q}$, respectively, while $t^{d}$ represents the target time of the desired image), the ReID system sends these information to each edge device. Then, for the $i$-th edge device, it extracts the deep feature $\mathbf{f}^{q}$ from $\mathcal{I}^{q}$ via a local visual backbone and compute the similarity $\mathbf{s}^{i}\in\mathbb{R}^{N_{i}}$ between $\mathbf{f}^{q}$ and the features $\mathbf{G}^{i}\in\mathbb{R}^{N_{i}\times E}$ of all $N_{i}$ gallery images on the $i$-th device ($E$ is the dimension). The images in current edge device are ranked by $\mathbf{s}^{i}$ and sent to the cloud server in batches due to the limited bandwidth. Finally, users check the returned data, and the process can be terminated if the desired images are contained in current batch.
Figure 2: The overview of our proposed cloud-edge collaborative inference framework. The DaCM is deployed in both the cloud server and edge devices for adjusting uploading batch size $b^{i}$ and image order in the uploading sequence. The red solid denotes the data flow enabled by the designed DaCM, and the gray dashed line denotes the previous data flow that can be removed by DaCM.
Denoting the user desired image as $\mathcal{I}^{d}$, we can see that the performance of cloud-edge collaborative ReID systems depends on when the $\mathcal{I}^{d}$ is returned to the server, which is influenced by: (1) the delay of the communication network; (2) the speed of feature extraction; (3) the rank of $\mathcal{I}^{d}$ in the uploading sequence, and (4) the amount of data that the camera (on which $\mathcal{I}^{d}$ is captured) can upload each time, i.e. the bandwidth allocated for the camera. The first two problems have been well studied by previous methods (Zhuang et al. 2020; Zhuang, Wen, and Zhang 2021; Jain et al. 2020; Kang et al. 2017; Zhang et al. 2020), but the last two problems still lack relevant methods. Therefore, in this paper, we focus on how to accelerate the system by advancing the position of $\mathcal{I}^{d}$ in the sequence and increasing the bandwidth utilization rate of the camera where $\mathcal{I}^{d}$ is located from the perspective of multimedia computing. The optimization objective to reduce the Transmission Number (TN) can be expressed as:
@@
where $\mathcal{U}^{i}$ is the uploading image sequence of $i$-th edge device, which is determined by the score $\mathbf{s}$, $b^{i}$ is the allocated bandwidth for $i$-th edge device, and it means how many images can be uploaded each time, $\theta$ is the parameters should be optimized, which influence $\mathcal{U}^{i}$ and $b^{i}$. Function $\epsilon$ is used to return the index of $\mathcal{I}^{d}$ in $\mathcal{U}^{i}$, function $\Omega$ aggregate the results from all edge devices, and their implementations are contingent upon user requirements.
Most of previous methods pay much attention to learning a proper $\mathcal{U}^{i}$, i.e. forcing the images with same identity ID $l^{q}$ have small $\mathbf{s}$ to make them at the front of the sequence, and they can not have an impact on $b^{i}$. In contrast, we propose the DaCM network, which can learn the spatial-temporal distribution of objects in scene and be used to boost the efficiency of the system by adjusting both $\mathcal{U}^{i}$ and $b^{i}$.
@@ 4 Proposed Approach
In this section, we first present the details of DaCM architecture and its training strategy, then describe how DaCM is used in the cloud-edge collaboration framework.
@@ 4.1 DaCM Architecture
The DaCM network performs an important role in the cloud-edge collaborative ReID system, and in this part, we describe its architecture. As shown in Fig. 3, DaCM mainly consists of three components, a spatial-temporal embedding module, multiple Correlation Modeling (CoMo) blocks, and a final classifier.
Spatial-temporal embedding. Inspired by the positional encoding manner used in (Vaswani et al. 2017), we first adopt the sinusoidal embedding to encode the temporal information. Denoted the timestamps of query image and target as $t^{q}$ and $t^{d}$, we encode their difference to a feature vector in the formulation of:
$\mathbf{e}_{2i}=\sin\left(\frac{(t^{d}-t^{q})}{\lambda^{\frac{2i}{D}}}\right),~{}\mathbf{e}_{2i+1}=\cos\left(\frac{(t^{d}-t^{q})}{\lambda^{\frac{2i}{D}}}\right),$ (2)
where $i$ is the dimension index, $\mathbf{e}\in\mathbb{R}^{D}$ is the results embedding, and $D$ is the dimension of the embedding. $\lambda$ denotes the max period of the sinusoidal function. The wavelengths form a geometric progression from $2\pi$ to $\lambda\cdot 2\pi$.
For the spatial information, considering the fixed topology of different cameras, we suggest employing learnable parameters $\mathbf{W}\in\mathbb{R}^{C\times D\times C\times D}$ to represent the spatial information. $C$ is the number of cameras and the nodes in the graph. When $c^{q}$ is provided as input, we choose $\mathbf{W}_{c^{q}}$ as the resulting spatial embedding, which is then combined with $\mathbf{e}$ to generate a graph structure:
$\mathbf{A}=\frac{1}{\sum_{j}^{D}e_{j}}\sum_{j}^{D}\mathbf{W}_{c^{q},j}*e_{j}+\mathbf{b}\in\mathbb{R}^{C\times D},$ (3)
where $\mathbf{b}$ is a learnable bias.
Correlation modeling block. As for CoMo blocks, they are used to propagate the information among graph nodes, and as shown in Fig. 3, one CoMo block is comprised of a Layer Normalization, a GCN layer (Kipf and Welling 2017), and a GeLU activation, which can be formulated as:
@@
where $\mathbf{E}^{l}\in\mathbb{R}^{C\times C}$ is the adjacency matrix of the GCN layer, which can be learned during training, and $\mathbf{W}^{l}\in\mathbb{R}^{C\times C}$ denotes the transfer weight to update the information for each node feature.
@@
The MLP consists of the sequence of BatchNorm (Ioffe and Szegedy 2015), ReLU, and fully connected layers.